OpenStudy (anonymous):

Can someone explain to me how to do this? Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

hero (hero):

This is very easy to do. All you have to do is take points (-4,0) and (4, y) and insert them into the formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) Upon doing so, you get: \((x - (-4))^2 + (y - 0)^2 = (x - 4)^2 + (y - y)^2\)

hero (hero):

Okay, now simplify that to get: \((x + 4)^2 + y^2 = (x - 4)^2\)

hero (hero):

Then expand the squares to get: \(x^2 + 8x + 16 + y^2 = x^2 - 8x + 16\)

hero (hero):

Notice that \(x^2\) and \(16\) cancel: \(8x + y^2 = -8x\)

OpenStudy (anonymous):

so the answer should be y^2=-16x?

hero (hero):

Next continue solving for x: \(y^2 = -16x\) \(-\dfrac{y^2}{16} = x\)

hero (hero):

Isolating \(x\) is the correct thing to do here.

hero (hero):

Sometimes \(x\) is expressed as \(x = -\dfrac{1}{16}y^2\)

OpenStudy (anonymous):

ok I got it thank you :)

hero (hero):

You can use this formula as long as you have a focus and a directrix.

hero (hero):

Just make sure you keep up with whether or not the directrix is equal to x or y.

hero (hero):

In this case, if you notice, the directrix was x = 4. The directrix, in this case, can also be expressed as (4, y)

hero (hero):

Do you know why that is?

OpenStudy (anonymous):

Because it says it's equal to x?

OpenStudy (anonymous):

I don't know exactly but it makes sense